OpenStudy (anonymous):

the ratio between two numbers is 12:13. If each number reduced by 20,the ratio becomes 2:3 .find the numbers.

OpenStudy (princeharryyy):

a/b = 12/13 (a-20)/(b-20) = 2/3 Solve both the equations for a and b